The following hashtable has a capacity of 10, and __ marks empty slots. The index of each slot is shown below for convenience. Using linear probing (probe length = 1), you insert a new key x (where x is your student id) into the hashtable. Next, you insert a second new key y (where y is your student id + 10). In which slot will y be inserted?(JAVA)

[ 60 __ __ 53 23 __ 36 46 27 __ ]

<0> <1> <2> <3> <4> <5> <6> <7> <8> <9>
Answer #1

Linear probing:
Linear probing is a form of open addressing for resolving collisions in hash table. Here, gap between two probes is 1.

If the key slot is full then we find the next empty slot and then insert key.

Key 60 __ __ 53 23 __ 36 46 27 __
Index 0 1 2 3 4 5 6 7 8 9

Hash function : hash(key) = key mod 10

Index : hash(key) % capacity

Inserting x :
Let's assume x = 71
So, hash(x) = 71 % 10 = 1.
So, the index = 1 % 10 = 1
since, index = 1 is empty so, we insert 71 at index 1.

Inserting y:
Given y = x + 10.
So the hash function will give the same value.
So, hash(x) = hash(x+10) = hash(y).
because, x%10 = (x+10)%10

So, y = 71 + 10 = 81 (since x = 71)
hash(y) = 81 % 10 = 1
index of y = 1 % 10 = 1
Since, index = 1 is occupied by x so we probe until we find the next empty slot.
next empty index is index = 2.

So, we insert y at index 2.

Key 60 x=71 y=81 53 23 __ 36 46 27 __
Index 0 1 2 3 4 5 6 7 8 9

If the answer helped please upvote, it means a lot and for any query please comment.